# A Survey on Wi-Fi Sensing Generalizability: Taxonomy, Techniques, Datasets, and Future Research Prospects

### v1

@@ 3. Dataset
    SHARP (Meneghello et al., 2022) dataset was collected in three environments: a bedroom, a living room, and a laboratory, featuring four activities—sit, walk, run, and jump—along with an empty room state. In the bedroom, volunteers p1 and p2 performed at two locations; in the living room, volunteer p1 performed at one location; and in the laboratory, volunteer p3 performed. These activities were conducted at different times over ten months. Each activity lasted 120 seconds per session, resulting in a total of 120 minutes of data. The setup included a single-antenna transmitter and receiver, along with a 4-antenna Asus RT-AC86U router as a monitoring device for collecting Wi-Fi data at a sampling rate of 173 Hz.
  40. (40)
@@
    XRFv2 (Lan et al., 2025) dataset is designed for continuous action localization and action summarization tasks. It was collected in three distinct indoor settings—a study room, a dining room, and a bedroom. In each environment, 16 volunteers performed continuous sequences of actions spanning 30 categories. The dataset comprises 853 continuous action sequences, each lasting approximately 80 seconds, making it suitable for continuous action detection, localization, and summarization tasks. The data collection setup utilized one transmitter and three receivers with a sampling rate of 200 Hz. In addition to Wi-Fi measurements, participants wore six IMU sensors placed on an earphone, glasses, two phone pockets, and two wrist-worn watches. Furthermore, XRFv2 provides pose annotations, supporting multimodal fusion and cross-modal research.
4. Challenges and Future Directions
4.1. Data, Data, Data
@@
The scaling law, which involves training large-scale models with large-scale datasets, has proven its success since the advent of AlexNet (Krizhevsky et al., 2012), achieving remarkable breakthroughs in fields such as computer vision (CV), natural language processing (NLP), robotics, and science. Large-scale datasets like ImageNet (Deng et al., 2009), COCO (Lin et al., 2014), Kinetics (Kay et al., 2017), and KITTI (Geiger et al., 2012) have significantly accelerated advancements in these areas. In contrast, the field of Wi-Fi human sensing has not progressed as rapidly, despite the widespread adoption of deep learning methods. A key reason lies in the limited scale and diversity of available Wi-Fi sensing datasets.
@@
Although researchers have made significant efforts to close this gap by collecting and releasing datasets like XRF55 (Wang et al., 2024b) and Widar 3.0 (Zheng et al., 2019), these datasets still suffer from constraints in terms of collection scenarios, action categories, and participant diversity. When compared to the scale of video, image, or text datasets, Wi-Fi sensing datasets lag behind by several orders of magnitude.
@@
  1. (1)
    How to model the propagation of signals through people and environments? Can methods like ray-tracing and Fresnel zone models effectively simulate Wi-Fi signals’ interactions with objects and people? How to enable modern approaches, such as diffusion models (Ho et al., 2020), Variational Autoencoders (VAEs) (Kingma and Welling, 2013), Neural Radiance Fields (NeRF) (Mildenhall et al., 2021), or 3D Gaussian Splatting (Kerbl et al., 2023) learn probabilistic mappings of signal propagation patterns from observed data?
  2. (2)
    How to construct diverse environments for data generation? Should we rely on manual modeling of environments, even though it is resource-intensive? Or should we leverage open-source tools like Unity, Unreal Engine, or Infinigen Indoors (Raistrick et al., 2024), which provide scalable solutions for generating dynamic, realistic environments tailored to Wi-Fi sensing scenarios?
  3. (3)
    How to synthesize diverse human body models? Is it more effective to use game engines like Unity or Unreal Engine to create diverse human models? Alternatively, could tools such as SMPL (Loper et al., 2015) or MANO (Romero et al., 2022) be utilized to extract human body models from large-scale video datasets for synthetic data generation?
  4. (4)
    How to bridge the gap between synthetic and real data? Could Generative Adversarial Networks (GANs) (Goodfellow et al., 2014) help enhance the realism of synthetic features or adapt synthetic data to better match real-world characteristics? What role could domain adaptation and domain-invariant feature learning play in aligning synthetic and real data to improve model performance?
@@
4.2. Wi-Fi Sensing Foundation Model Pre-Training
Foundation models in vision and language, such as CLIP (Radford et al., 2021), GPT (Brown et al., 2020), and BERT (Devlin et al., 2019), have demonstrated remarkable success in adapting to diverse downstream tasks. Similarly, Wi-Fi sensing encompasses a wide range of applications, including human activity recognition, indoor localization, presence detection, and pose estimation. These varied tasks make the development of a Wi-Fi sensing foundation model both promising and necessary. However, training such a model raises several critical open questions that must be addressed:
  1. (1)
    Should we utilize traditional architectures like CNNs and RNNs, or opt for Transformers (Waswani et al., 2017), which have demonstrated strong performance across various domains? Alternatively, emerging architectures such as Mamba (Gu and Dao, 2023), known for their smaller parameter sizes, reduced memory consumption, and faster inference speeds, may also be viable options. Should we further tailor the architecture design to cater to different computational platforms, optimizing for specific hardware constraints and efficiency requirements?
  2. (2)
@@ 4.2. Wi-Fi Sensing Foundation Model Pre-Training
  3. (3)
    What self-supervised learning strategy is most suitable? Self-supervised learning can effectively leverage unlabeled data, but how should we design the proxy tasks? Should we employ contrastive learning strategies that bring positive sample pairs closer while pushing negative pairs apart? Or would a masking-then-reconstruction approach, as used in models like BERT (Devlin et al., 2019) or MAE (He et al., 2022), be more appropriate for Wi-Fi sensing? Additionally, are there domain-specific proxy tasks that could better capture the unique spatiotemporal characteristics of Wi-Fi signals?
  4. (4)
@@
4.3. When Wi-Fi Sensing Meets Large Multimodal Models
Beyond training Wi-Fi foundation models from scratch, as discussed in Section 4.2, another promising research direction is leveraging existing large multimodal models by fine-tuning them to process Wi-Fi sensing data. Large multimodal models, such as those trained on diverse combinations of images, text, and audio, have demonstrated impressive adaptability across various domains. By aligning Wi-Fi sensing data with these models, researchers may unlock new possibilities for enhanced understanding and application of Wi-Fi sensing technologies. However, integrating Wi-Fi data into such models raises several open questions and challenges:
  1. (1)
    What fine-tuning strategy is most effective? Fine-tuning large multimodal models requires balancing computational efficiency with task-specific performance improvements. Techniques like LoRA (Hu et al., 2022), adapters (Rebuffi et al., 2017), or prompt-based fine-tuning (Liu et al., 2021) are popular in the NLP and vision domains. Should one of these strategies be adopted for Wi-Fi sensing, or is there a need for a novel approach tailored to the unique characteristics of Wi-Fi data? Moreover, how can we efficiently fine-tune large models with limited computational resources while ensuring high adaptability?
  2. (2)

### v2

@@ 42 CSI-Bench [73] 26 environments, 35 participants 16 device configuration In-the-wild data, 461 hours
(17) SHARP [91] dataset was collected in three environments: a bedroom, a living room, and a laboratory, featuring four activities—sit, walk, run, and jump—along with an empty room state. In the bedroom, volunteers P1 and P2 performed at two locations; in the living room, volunteer p1 performed at one location; and in the laboratory, volunteer P3 performed. These activities were conducted at different times over ten months. Each activity lasted 120 seconds per session, resulting in a total of 120 minutes of data. The setup included a single-antenna transmitter and receiver, along with a 4-antenna Asus RT-AC86U router as a monitoring device for collecting Wi-Fi data at a sampling rate of 173 Hz.
(18) CSI-HAR-Dataset [291] dataset was collected in a single indoor environment. The dataset was collected by 3 volunteers of different ages (adult, middle-aged, and elderly). Each volunteer performed seven categories of actions: walking, running, falling, lying down, sitting down, standing up, and bending. Each action was repeated 20 times by each participant, resulting in 420 samples. The experimental setup involved a Raspberry Pi 4 as the receiver and a Tp-link Archer C20 wireless router as the transmitter, operating on a 5 GHz WiFi network with a 20 MHz bandwidth on channel 36 (IEEE 802.11ac standard). The transmitter and receiver were positioned 3 meters apart, both 1 meter above the ground to ensure an unobstructed signal path. The sampling rate was 200 Hz, with 4000 samples collected over 20 seconds per activity, where the activity occurred in the middle of this period (approximately 3–6 seconds).
(19) CSIDA, related in WiGr [193], was collected in an office environment with three distinct locations and a classroom environment with two locations. Five volunteers performed six gestures: pull left, pull right, lift up, press down, draw a circle, and draw zigzag. Each gesture was repeated 10 times, with each repetition lasting 1.8 seconds. The device setup consisted of a single-antenna transmitter and a three-antenna receiver, separated by a distance of 2.6 meters. The system operated in the 5 GHz mode with a 40 MHz channel bandwidth, capturing 114 subcarriers with a sampling rate of 1000 Hz.
(20) RISE [237] dataset includes data for gesture recognition and gait recognition, collected in two environments: an office and a controlled setting with a radio frequency anechoic chamber to minimize multipath effects. Gesture Recognition: Six volunteers performed six gestures—push and pull, draw a circle, throw, slide, sweep, and draw zigzag—across ten locations (five per environment). Each gesture was repeated 30 times per location, resulting in 10,800 samples (6 volunteers ×6 gestures ×10 locations ×30 repetitions). Gait Recognition: Fifteen volunteers walked under five different configurations in the controlled environment, generating 1,150 samples across ten activity types.
(21) HTHI [292] dataset was collected in a furnished room. 66 volunteers participated in the study, forming 40 unique pairs to perform 12 types of human-to-human interaction activities, such as handshaking, high fives, and hugging. Each pair performed each activity 10 times, resulting in 4,800 samples. The data collection setup consisted of a commercial off-the-shelf Sagemcom 2704 router with two antennas acting as the transmitter and a desktop computer with three antennas serving as the receiver. The transmitter and receiver were positioned 4.3 meters apart. The sampling rate for the data collection was not reported.
(22) DeepSeg  [261] dataset was collected in a meeting room with data from 5 volunteers performing 10 activities. These actions include 5 fine-grained movements such as swinging the hand, raising the hand, making a pushing motion, tracing a circular pattern (O), and tracing a cross pattern (X), as well as 5 coarse-grained activities, namely boxing, picking up an object, running, squatting down, and walking. Each participant repeated each action 30 times, resulting in 1,500 samples. The dataset collection setup included a single-antenna transmitter and a three-antenna receiver placed 2 meters apart, with a sampling rate of 50 Hz.
(23) RF-Net [199] dataset was collected in six different rooms with 11 volunteers performing six types of actions: wiping, walking, moving, rotating, sitting, and standing up. The Wi-Fi setup included two transmitter-receiver pairs, operating at a sampling rate of 100 Hz.
dataset was collected in six different rooms with 11 volunteers performing six types of actions: wiping, walking, moving, rotating, sitting, and standing up. The Wi-Fi setup included two transmitter-receiver pairs, operating at a bandwidth of 20 MHz and a sampling rate of 100 Hz. Each participant performed each action 20 times in 100 environments, resulting in 12,000 samples.
(24) Baha’A et al. [293] data collected in three different environments: two with line-of-sight (LOS) settings and one with a non-line-of-sight (NLOS) setting. Thirty volunteers performed five different actions: falling from a sitting position, falling from a standing position, walking, sitting down and standing up, and picking up a pen from the ground. Each action was performed 20 times. The experimental setup included a single-antenna transmitter and a three-antenna receiver, operating at a frequency of 2.4 GHz with a sampling rate of 320 Hz.
(25) Widar3.0 [67] dataset was collected in three environments: a classroom, a spacious hall, and an office room. In each environment, volunteers performed actions within a 2m $\times$ 2m area, with five designated locations and five orientations at each location. The dataset includes two types of hand gestures. The first type consists of general hand movements, including pushing and pulling, sweeping, clapping, sliding, drawing circles, and drawing zigzags. The second type involves drawing the numbers 0$\sim$9. The first type of action was performed by 16 participants, resulting in 12,000 samples (16$\times$5$\times$5$\times$6$\times$5 instances), while the second type was demonstrated by 2 participants, with 5,000 samples (2$\times$5$\times$5$\times$10$\times$10 instances). Widar3.0 was collected using a transmitter with one antenna and at least three receivers, each with three antennas. The dataset was captured on channel 165 at 5.825 GHz with a sampling rate of 1,000 Hz.
(26) ARIL [15] dataset was collected in a 3.2m $\times$ 3.2m area within a laboratory, with 16 positions. The actions include 6 hand gestures for potential human-computer interaction applications: hand up, hand down, hand left, hand right, hand circle, and hand cross. One volunteer performed each gesture 15 times at each position. The dataset contains a total of 1,394 valid samples. Data was collected using two USRPx210 devices, each with a single antenna, as transmitter and receiver. The sampling rate was approximately 60 Hz.
(27) WiAR [294] dataset was collected in three indoor environments: one empty room, one meeting room, and one office, as well as two outdoor environments, including a playground and a corridor. Ten volunteers performed three categories of actions: upper body activities, lower body activities, and whole-body activities, with 16 different actions. Each action was performed 30 times. The experimental setup included a transmitter with a single antenna and one receiver with three antennas, spaced 4 meters apart. The participants performed the actions at the center of the setup.
(28) SAR  [295, 221] dataset was collected in the living room of an apartment. A total of 9 volunteers participated in experiments over 6 days. During the first 3 days, 3 different volunteers participated each day, while 2 volunteers participated repeatedly over the last 3 days. During the first 3 days, participants performed 6 actions: Clap, Walk, Wave, Jump, Sit, and Fall. In the last 3 days, participants performed the same actions except for Jump, resulting in a total of 5 actions. Each action was performed continuously for 5 seconds and repeated 50 times. The data collection setup consisted of a transmitter and a receiver, each equipped with three antennas, spaced approximately 2.5 meters apart. The sampling rate was 20 Hz.
(29) FallDeFi [39], designed for fall detection, was collected in five environments: a kitchen, bedroom, corridor, laboratory, and bathroom. Three participants performed fall actions and other activities such as walking, jumping, sitting down, and standing up, with each action lasting for 10 seconds and 1153 samples in total. Each environment featured one transmitter and one receiver, with varying placement distances between 4 and 9 meters depending on the setting. To introduce environmental variations, data was collected on different days, and the transmitter’s position was adjusted during the experiments. All experiments were conducted at 5.2 GHz with a sampling rate of 1000 Hz and a bandwidth of 20 MHz.
(30) SignFi [14] dataset was collected in two environments: a laboratory and a home. The lab, measuring 13m $\times$ 12m, the distance between the AP and the STA was 230 centimeters. While the home environment, with dimensions of 4.11m $\times$ 3.86m, offered a shorter distance of 130 centimeters between the AP and STA. 5,520 instances were recorded in the lab, while the home environment provided 2,760 instances. The data involved 276 different sign language gestures, encompassing movements of the head, arm, hand, and fingers. The gestures were performed by 5 different users, and the users were instructed to perform each sign gesture repeatedly, with 20 instances for the lab environment and 10 instances for the home environment, resulting in a total of 8280 gesture instances. The experiments were conducted at 5 GHz with a bandwidth of 20 MHz.
(31) CrossSense [158] dataset was collected in three scenarios: a hall entrance, a narrow corridor, and a room, with the participation of 100 volunteers. For gait recognition, each volunteer walked 20 times in the same direction in each scenario, resulting in a total of 6,000 samples. For action recognition, five predefined positions were set in each scenario. At each position, each volunteer performed 40 different actions (e.g., pull, kick, throw), repeating each action 10 times. This resulted in a total of 600,000 samples. The experimental setup included one transmitter and one receiver, both equipped with three antennas. The devices operated at 5 GHz with a sampling rate of 1,000 Hz.
(32) Yousef et.al. 2017 [62]dataset was collected in an indoor office area where the WiFi transmitter and receiver were positioned 3m apart in a line-of-sight (LOS) condition. Six volunteers performed six different activities: lying down, falling down, walking, running, sitting down, and standing up. Each activity was performed 20 times, yielding 720 samples, and each activity was recorded for 20 seconds. Additionally, video recordings were made during the experiments to help label and annotate the activities in the dataset. The sampling rate was 1000 Hz.
(33) NTU-Fi HumanID [191] dataset was collected in a laboratory and a cubic office. Twenty volunteers, including 12 males and 8 females aged between 20 and 28, participated in the experiments, with 15 designated as legal users and the remaining as illegal intruders. Volunteers walked individually through the testing area, and their gait was captured using WiFi signals. Two TP-Link N750 routers were used: one as a transmitter with a single antenna, operating in 802.11n AP mode at 5 GHz with a 40 MHz bandwidth, and the other as a receiver with three antennas.
(34) GaitID [30] dataset is designed for gait recognition tasks. Data was collected in two scenarios: a hall and a discussion room. Volunteers walked along four predefined tracks, each with two walking directions (starting from both ends of each track). In the hall, 10 volunteers participated, with each walking 50 times in each direction. In the discussion room, 3 volunteers repeated each track 25 times in each direction. A total of 4,600 samples were collected. The data collection setup included one transmitter and six receivers, forming six Wi-Fi links. The sampling rate, bandwidth, and center frequency were not reported.
@@
(35) XRFv2 [296] dataset is designed for continuous action localization and action summarization. It was collected in three distinct indoor settings—a study room, a dining room, and a bedroom. In each environment, 16 volunteers performed continuous actions spanning 30 categories. The dataset comprises 853 continuous action sequences, each lasting approximately 80 seconds, making it suitable for continuous action detection, localization, and summarization tasks. The data collection setup utilized one transmitter and three receivers with a sampling rate of 200 Hz. In addition to Wi-Fi measurements, participants wore six IMU sensors placed on an earphone, glasses, two phone pockets, and two wrist-worn watches. Furthermore, XRFv2 provides pose annotations, supporting multimodal fusion and cross-modal research.
(36) Person-in-WiFi 3D [50] dataset was collected in three environments: an office, a classroom, and a corridor, each within a 4m $\times$ 3.5m area. Seven participants freely performed eight actions: reaching out, raising hands, bending over, stretching, sitting down, lifting legs, standing, and walking. In addition to scenarios with a single participant, there were also cases where 2, 3, or 4 participants simultaneously performed actions within the scene. Each clip was recorded for 40 seconds, resulting in a total of 456 40-second recordings. The data collection setup included one transmitter with a single antenna and three receivers, each with three antennas, arranged in a quadrilateral layout. The sampling rate was 200 Hz.
(37) MetaLoc [197] dataset was designed for indoor localization and collected in a 12 m × 5 m hall and a 10 m × 8 m laboratory. In each environment, Wi-Fi fingerprints (RSS and CSI) were recorded at 90 grid points. The setup included three transmitters: an ASUS RT-AC86U, a TP-Link TL-WR885N, and a TP-Link TL-WR886N, and a Nexus-5 smartphone as the receiver. The system operated at 5 GHz with a bandwidth of 20 MHz.
(38) Chen and Chang [297] released a dataset used for WiFi fingerprinting localization tasks. Data was collected in two different environments: a conference room and a cubicle office. In the conference room, data was collected at 16 locations, with 600 samples recorded at each location. In the cubicle office, data was collected at 18 locations, with 500 samples recorded at each location. The data acquisition system in the conference room consisted of one fixed-location transmitter-receiver pair, each equipped with two antennas, with a distance of 6 meters between them. In the cubicle office, the data acquisition system consisted of two fixed-location transmitter-receiver pairs, each equipped with two antennas. The number of volunteers participating in data collection, sampling rate, bandwidth, and center frequency were not reported.
(39) The SDP dataset [298] is an industrial-grade Wi-Fi sensing dataset tailored for person detection and counting tasks, featuring realistic deployments in both home and office scenarios. In the home setting, two scenarios are included, each consisting of three rooms—one parlor equipped with an access point (AP) and two bedrooms with replay devices. Each device is equipped with two antennas and operates at 5GHz using the 802.11ax protocol, with a sampling rate of 20Hz and bandwidths of 80MHz or 160MHz. Up to two people may be present and moving in each room, and a sweeping robot is introduced as an additional source of interference. The office setting includes two scenarios: the first features four rooms, each with an AP and four-antenna devices, operating on the 5GHz band with 802.11ac and 20MHz bandwidth; the second involves a corridor with three APs placed at the left, center, and right, and four adjacent rooms without APs. This scenario captures up to two occupants per room and includes varying sampling rates—5Hz, 20Hz, 30Hz, 50Hz, and 100Hz—to explore the effect of temporal resolution.
(40) DASECount  [227] dataset, designed for crowd counting, was collected in two environments: an office and a lecture hall, each with both line-of-sight (LOS) and non-line-of-sight (NLOS) configurations, resulting in four different environmental settings. Volunteers, ranging from 0 to 8 people, performed three types of motion: static (seated with free actions like eating or typing), dynamic (random walking), and mixed (unrestricted activities including walking and sitting). Each sampling sequence was recorded for 5 minutes. The experimental setup included a two-antenna transmitter and a three-antenna receiver, operating at 2.4 GHz with a bandwidth of 40 MHz and a sampling rate of 100 Hz.
(41) WiFiTAD [299] is a dataset specifically designed for temporal action detection. It was collected in a single office with size of $7m\times 12m\times 2.5m$, where participants continuously performed a sequence of seven actions between the transmitter and receiver, including walking, running, jumping, waving, falling, sitting, and standing. The dataset comprises a total of 553 action sequences, each with an average duration of 85 seconds. Data were collected from 3 volunteers using a laptop equipped with an Intel 5300 NIC, configured with one transmitter and one receiver, each having a single antenna, with a sampling rate of 100Hz.
(42) CSI-Bench [73] is a comprehensive, in-the-wild Wi-Fi sensing benchmark that integrates multiple classification-oriented sub-datasets to support a wide range of sensing tasks. These include fall detection (6 environments, 17 participants, 6,700 samples), breathing detection (3 environments, 3 users, 100,000 samples), motion source recognition (10 environments involving 13 humans, 20 pets, a robot, and a fan), room-level localization (6 environments, 8 users), user identification (6 users), activity recognition (5 activity classes), and proximity recognition (4 distance levels). A key strength of CSI-Bench lies in its emphasis on realism: data are collected using commercial Wi-Fi edge devices across 26 diverse indoor environments with 35 real users. Spanning over 461 hours of effective recordings, CSI-Bench captures rich signal variations under natural, uncontrolled conditions, making it a valuable resource for evaluating the generalization, robustness, and scalability of Wi-Fi sensing algorithms.
(43) CSI-Net [210] is a multi-task Wi-Fi sensing dataset collected in a single indoor environment ($5m\times 6m$) using one transmitter and one receiver, each with three antennas, and a sampling rate of 100Hz. It includes four sub-tasks: (1&2) User Identification and Biometrics Estimation, involving 30 participants who remained stationary for 100 seconds while their CSI signals were recorded at a distance of 1.6m; additional biometric attributes such as body fat and muscle rate were also provided; (3) Sign Language Recognition, in which one participant performed 10 American Sign Language digits (0–9), each lasting about 60 seconds, with the transmitter-receiver distance set to 0.6m; and (4) Fall Detection, where the same participant simulated falls at five different locations in the room, with each fall motion lasting 30 seconds and a 3.0m separation between transmitter and receiver.
IV Challenges and Future Directions
IV-A Data, Data, Data
@@
The scaling law, which involves training large-scale models with large-scale datasets, has proven its success since the advent of AlexNet [300], achieving remarkable breakthroughs in fields such as computer vision (CV), natural language processing (NLP), robotics, and science. Large-scale datasets like ImageNet [301], COCO [302], and Kinetics [303] have significantly accelerated advancements in these areas. In contrast, the field of Wi-Fi human sensing has not progressed as rapidly, despite the widespread adoption of deep learning methods. A key reason lies in the limited scale and diversity of available Wi-Fi sensing datasets.
@@
Although researchers have made significant efforts to close this gap by collecting and releasing datasets like XRF55 [70] and Widar 3.0 [67], these datasets still suffer from constraints in terms of collection scenarios, action categories, and participant diversity. When compared to the scale of video, image, or text datasets, Wi-Fi sensing datasets lag behind by several orders of magnitude.
@@
  1. 1.
    How to model the propagation of signals through people and environments? Can methods like ray-tracing and Fresnel zone models effectively simulate Wi-Fi signals’ interactions with objects and people? How to enable modern approaches, such as diffusion models [262], Variational Autoencoders (VAEs) [304], Neural Radiance Fields (NeRF) [305], or 3D Gaussian Splatting [306] learn probabilistic mappings of signal propagation patterns from observed data?
  2. 2.
    How to construct diverse environments for data generation? Should we rely on manual modeling of environments, even though it is resource-intensive? Or should we leverage open-source tools like Unity, Unreal Engine, or Infinigen Indoors [307], which provide scalable solutions for generating dynamic, realistic environments tailored to Wi-Fi sensing scenarios?
  3. 3.
    How to synthesize diverse human body models? Is it more effective to use game engines like Unity or Unreal Engine to create diverse human models? Alternatively, could tools such as SMPL [308] or MANO [309] be utilized to extract human body models from large-scale video datasets for synthetic data generation?
  4. 4.
    How to bridge the gap between synthetic and real data? Could Generative Adversarial Networks (GANs) [252] help enhance the realism of synthetic features or adapt synthetic data to better match real-world characteristics? What role could domain adaptation and domain-invariant feature learning play in aligning synthetic and real data to improve model performance?
@@
IV-B Wi-Fi Sensing Foundation Model Pre-Training
Foundation models in vision and language, such as CLIP [310], GPT [311], and BERT [255], have demonstrated remarkable success in adapting to diverse downstream tasks. Similarly, Wi-Fi sensing encompasses a wide range of applications, including human activity recognition, indoor localization, presence detection, and pose estimation. These varied tasks make the development of a Wi-Fi sensing foundation model both promising and necessary. However, training such a model raises several critical open questions that must be addressed:
  1. 1.
    Should we utilize traditional architectures like CNNs and RNNs, or opt for Transformers [312], which have demonstrated strong performance across various domains? Alternatively, emerging architectures such as Mamba [313], known for their smaller parameter sizes, reduced memory consumption, and faster inference speeds, may also be viable options. Should we further tailor the architecture design to cater to different computational platforms, optimizing for specific hardware constraints and efficiency requirements?
  2. 2.
@@ IV-B Wi-Fi Sensing Foundation Model Pre-Training
  3. 3.
    What self-supervised learning strategy is most suitable? Self-supervised learning can effectively leverage unlabeled data, but how should we design the proxy tasks? Should we employ contrastive learning strategies that bring positive sample pairs closer while pushing negative pairs apart? Or would a masking-then-reconstruction approach, as used in models like BERT [255] or MAE [314], be more appropriate for Wi-Fi sensing? Additionally, are there domain-specific proxy tasks that could better capture the unique spatiotemporal characteristics of Wi-Fi signals?
  4. 4.
@@
IV-C When Wi-Fi Sensing Meets Large Multimodal Models
Beyond training Wi-Fi foundation models from scratch, as discussed in Section IV-B, another promising research direction is leveraging existing large multimodal models by fine-tuning them to process Wi-Fi sensing data. Large multimodal models, such as those trained on diverse combinations of images, text, and audio, have demonstrated impressive adaptability across various domains. By aligning Wi-Fi sensing data with these models, researchers may unlock new possibilities for enhanced understanding and application of Wi-Fi sensing technologies. However, integrating Wi-Fi data into such models raises several open questions and challenges:
  1. 1.
    What fine-tuning strategy is most effective? Fine-tuning large multimodal models requires balancing computational efficiency with task-specific performance improvements. Techniques like LoRA [315], adapters [316], or prompt-based fine-tuning [317] are popular in the NLP and vision domains. Should one of these strategies be adopted for Wi-Fi sensing, or is there a need for a novel approach tailored to the unique characteristics of Wi-Fi data? Moreover, how can we efficiently fine-tune large models with limited computational resources while ensuring high adaptability?
  2. 2.